# Automated Contrastive Learning Strategy Search for Time Series

### v1

@@ 5. Conclusion and Future Work
In this study, we introduced an innovative framework, AutoCL, for the automated discovery of Contrastive Learning Strategies (CLS) in time series data, aimed at diminishing the reliance on extensive domain expertise and iterative experimentation in the development of CLS tailored to specific datasets and tasks. The AutoCL framework is structured around two principal components: a comprehensive solution space, meticulously designed to encapsulate four critical dimensions of CLS, and an effective optimization algorithm to identify suitable and transferable CLS configurations. Our empirical tests on eight datasets demonstrate AutoCL’s effectiveness in finding suitable CLS, with the derived Generally Good Strategy (GGS) showing robust performance across different tasks. This study not only showcases the potential of AutoCL but also provides valuable insights for future contrastive learning research.
Given the versatility of our automation approach, we are confident that our proposed framework can be adapted to non-contrastive self-supervised learning approaches involving time series data, an avenue we intend to explore in our future research.
@@ Appendix A Data Augmentations
Data augmentations transforms the input data into different but related views, which are the cornerstones of a CLS. In general, each augmentation is associated with a parameter $p\in[0,1]$. Let $\mathbf{x}\in\mathbb{R}^{T\times c}$ be the input time series, where $T$ and $c$ are the length and the number of variables, then the details of data augmentations are given below.
@@ 6 Crop Point Mask Point Mask Point Mask Point Mask
Appendix B Procedure of Composing GGS
In this section, we elaborate the details for the procedure of composing GGS.
First, collect top 10 candidate CLS sets $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, $\mathcal{A}_{3}$ from HAR, ETTh1 and Yahoo, and construct candidate pairs: $(A_{1,i},A_{2,j},A_{3,k})$, where $A_{1,i}\in\mathcal{A}_{1}$, $A_{2,j}\in\mathcal{A}_{2}$, $A_{3,k}\in\mathcal{A}_{3}$.
Second, rank the candidate pairs by the number of their shared options. Note that for a probability option (e.g., jittering) $p\in[0.0,0.1,\cdots,0.9,0.95]$, it will be considered as a shared option, if $(\max(p_{1},p_{2},p_{3})-\min(p_{1},p_{2},p_{3}))<0.35$, where $p_{1}$, $p_{2}$, $p_{3}$ correspond to $A_{1,i},A_{2,j},A_{3,k}$.
Third, pick the pair that has the most shared options, and add their shared options in GGS. For the rest un-shared options, we search their values based on heuristics. The strategies with significant validation performance drop on any of HAR/ETTh1/Yahoo are discarded. We stop searching when a strategy with acceptable validation performance is found.
Appendix C More Experimental Setups
C.1. Evaluation Metrics
Our evaluation follows (Yue et al., 2022). For each dataset, we first pre-train an encoder based on a CLS, and then train a downstream model, which is evaluated on the test data. For classification, the downstream model is the SVM classifier with RBF kernel, and the evaluation metrics are accuracy (ACC) and F1. For forecasting, the linear regression model with $l_{2}$ norm is used, and the evaluation metrics are Mean Squared Error (MSE) and Mean Absolute Error (MAE). For anomaly detection, we follow the protocol used by (Yue et al., 2022; Ren et al., 2019). The evaluation metrics are F1, precision and recall.
C.2. Baselines
We consider the following baselines: InfoTS (Luo et al., 2023) adaptively learns optimal augmentations. TS2Vec (Yue et al., 2022) performs contrastive learning in an hierarchical manner. TS-TCC (Eldele et al., 2021) uses temporal and contextual contrasting to train encoders. CoST (Woo et al., 2021) performs augmentations in both temporal and frequency domains. TNC (Tonekaboni et al., 2020) treats adjacent samples as positive pairs and non-adjacent samples as negative pairs. CPC (Oord et al., 2018) introduces InfoNCE. Self-EEG (Sarkar and Etemad, 2020) uses various transformations to augment the input time series. SR (Ren et al., 2019) detects anomaly points based on spectral residual. DONUT (Xu et al., 2018) detect anomalies based on Variational Auto-Encoder (VAE) (Doersch, 2016). SPOT and DSPOT (Siffer et al., 2017) use the extreme value theory (Beirlant et al., 2006) to find anomalies. Note that GGS (Section 3.3) is the generally good strategy found by AutoCL, which has strong performance on different datasets and tasks.
C.3. Implementation Details
For the controller network, its hidden dimension size is 320, and MLPs are linear layers with a softmax activation. We use a 10 layer dilated CNN as the encoder similar to (Yue et al., 2022), and the sizes of hidden dimension and the output embeddings are 64 and 320. The learning rate of the controller is 0.0001. We fix the maximum number of the iterations in phase 1 as 500, and $\alpha$ as 10. For classification, forecasting and anomaly detection, the tolerance $\epsilon$ are 0.001, 0.0001 and 0.001, and the validation metrics are ACC, MSE (48 horizon) and F1. In pre-training, we fix the maximum length of the input time series as 2,000.

### v2

@@ 4. Conclusion and Future Work
In this study, we introduced an innovative framework, AutoCL, for the automated discovery of Contrastive Learning Strategies (CLS) for time series, aimed at diminishing the reliance on extensive domain expertise and iterative experimentation in the development of CLS tailored to specific datasets and tasks. AutoCL is structured around two principal components: a comprehensive solution space, meticulously designed to encapsulate four critical dimensions of CLS, and an effective optimization algorithm to identify suitable and transferable CLS configurations. Empirical evaluation on eight datasets demonstrates AutoCL’s effectiveness in finding suitable CLS, with the derived Generally Good Strategy (GGS) showing robust performance across different tasks. This study not only showcases the potential of AutoCL but also provides valuable insights for future contrastive learning research.
@@ Appendix A Data Augmentations
Data augmentations transform the input data into different but related views, which are the cornerstones of a CLS. In general, each augmentation is associated with a parameter $p\in[0,1]$. Let $\mathbf{x}\in\mathbb{R}^{T\times c}$ be the input time series, where $T$ and $c$ are the length and the number of variables, then the details of data augmentations are given below.
@@ 6 Crop Point Mask Point Mask Point Mask Point Mask
Appendix B More Experimental Setups
B.1. Evaluation Metrics
Our evaluation follows (Yue et al., 2022). For each dataset, we first pre-train an encoder based on a CLS, and then train a downstream model, which is evaluated on the test data. For classification, the downstream model is the SVM classifier with RBF kernel, and the evaluation metrics are accuracy (ACC) and F1. For forecasting, the linear regression model with $l_{2}$ norm is used, and the evaluation metrics are Mean Squared Error (MSE) and Mean Absolute Error (MAE). For anomaly detection, we follow the protocol used by (Yue et al., 2022; Ren et al., 2019). The evaluation metrics are F1, precision, and recall.
B.2. Baselines
We consider the following baselines: InfoTS (Luo et al., 2023) adaptively learns optimal augmentations. TS2Vec (Yue et al., 2022) performs contrastive learning in an hierarchical manner. TS-TCC (Eldele et al., 2021) uses temporal and contextual contrasting to train encoders. CoST (Woo et al., 2021) performs augmentations in both temporal and frequency domains. TNC (Tonekaboni et al., 2020) treats adjacent samples as positive pairs and non-adjacent samples as negative pairs. CPC (Oord et al., 2018) introduces InfoNCE. Self-EEG (Sarkar and Etemad, 2020) uses various transformations to augment the input time series. SR (Ren et al., 2019) detects anomaly points based on the spectral residual. DONUT (Xu et al., 2018) detect anomalies based on Variational Auto-Encoder (VAE) (Doersch, 2016). SPOT and DSPOT (Siffer et al., 2017) use the extreme value theory (Beirlant et al., 2006) to find anomalies. Note that GGS (Section 3.3) is the generally good strategy found by AutoCL, which has strong performance on different datasets and tasks.
B.3. Implementation Details
For the controller network, its hidden dimension size is 320, and MLPs are linear layers with a softmax activation. We use a 10 layer dilated CNN as the encoder similar to (Yue et al., 2022), and the sizes of hidden dimension and the output embeddings are 64 and 320. The learning rate of the controller is 0.0001. We fix the maximum number of the iterations in phase 1 as 500, and $\alpha$ as 10. For classification, forecasting, and anomaly detection, the tolerance $\epsilon$ are 0.001, 0.0001, and 0.001, and the validation metrics are ACC, MSE (48 horizons), and F1. In pre-training, we fix the maximum length of the input time series as 2,000.